Detemine whether the following graph can represent a variable with a normal distribution. Explain your reasoning. If the graph appears to represent a normal dstribution, estimate the mean and standard deviaton. 10 15 20 25 30 35 Could the graph represent a variable with a normal distribution? Explain your reasoning Select the correct choice below and, if necessary, fit in the answer boxes within your choice O A. No, because the graph is skewed let OB. Yes,the graph fulfis the properties of the nomal distribution. The mean is approximately and the standard deviation is about Type whole numbers) No, because the graph crosses the x-axis C, O D. No, because the graph is skewed right
Answer #1

After having a look at the graph we can assume that the graph represents a variable with a normal distribution.

Now from the graph it is clear that the mean is approximately 20 and the standard deviation can be find using the 3-\sigma limits.

From the graph we can see that 3*\sigma=(20-10)=(30-20)=10.

Hence \sigma =10/3=3.33

Hence the answer will be:

B.)Yes the graph fullfills the properties of the normal distribution. The mean is approximately 20 and the standard deviation is about 3.33

> bruh what

Jaden Cerlig Tue, Nov 16, 2021 4:34 PM
Answer #2

its 22 and 4

answered by: anonymous